OpenStudy (annie96):

I really need help. 1. 4 – 2y = 2 (9 – y) 2. 5 ( 3 – m) + 5m = 15 3. -2 |x| - 7 = -3

OpenStudy (anonymous):

are you solving for the variables?

OpenStudy (annie96):

Yes.

OpenStudy (anonymous):

3 is x=6

OpenStudy (anonymous):

Okay, well I am trying to solve the first one and the y's cancled, sooooo yeah..... idk what to do there

OpenStudy (the_fizicx99):

2*9 2*(-y) 4 – 2y = 18 -2y -4 -4 -2y = 14 - 2y +2y +2y y = 14 5 ( 3 – m) + 5m = 15 15 -5m + 5m = 15 15 = 15 so yes :P

OpenStudy (ranga):

4 – 2y = 2 (9 – y) 4 - 2y = 18 - 2y The 2y cancels out on both sides leaving with 4 = 18 which can never be true. Therefore, there is no solution for this equation. There is no value of y that can make 4 equal 18.

OpenStudy (the_fizicx99):

Oh thats what you're going for >.>

OpenStudy (ranga):

5 ( 3 – m) + 5m = 15 15 - 5m + 5m = 15 the 5m cancels out leaving with 15 = 15 which is always true. So all real numbers are solutions. So m is any real number.

OpenStudy (annie96):

Ok.

OpenStudy (ranga):

-2 |x| - 7 = -3 add 7 to both sides: -2|x| = 4 divide both sides by -2: |x} = -2 The absolute value of x can NEVER be negative. Therefore there is NO solution for this equation.

OpenStudy (ranga):

5th line above is supposed to be |x| = -2

OpenStudy (annie96):

Alright thank you.

OpenStudy (ranga):

You are welcome.